OpenStudy (anonymous):

Having trouble understanding these steps for one of my classes:

OpenStudy (anonymous):

\[zl = \theta*z(h-l)-\theta*G\] to \[zl*(1+ \theta) = \theta*zh-\theta*G\] to \[l = \theta/(1+\theta)*(h-G/z)\]

OpenStudy (anonymous):

Where all are constants except l

OpenStudy (anonymous):

Okay I get that, thanks a lot. Can you explain the last part too for us?

OpenStudy (jdoe0001):

hmmm what part again?

OpenStudy (anonymous):

From zl(1+Theta) = Thetazh -ThetaG to the answer where l is solved for

OpenStudy (anonymous):

l=θ/(1+θ)∗(h−G/z) is the right answer

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

sec

OpenStudy (anonymous):

No problem appreciate the help friend

OpenStudy (jdoe0001):

hmmm hmm lemme start off you want \(\bf zl = \theta*z(h-l)-\theta*G\qquad to\qquad l = \cfrac{\theta}{1+\theta}\cdot \cfrac{h-G}{z}\) right?

OpenStudy (anonymous):

yes